OpenStudy (anonymous):

Can anyone help me on algebra 2

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Thank you. im typing the question now

OpenStudy (calculusfunctions):

What do you need?

OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3

OpenStudy (anonymous):

f(x) is denoted by y. we are given the value of y as 4(x+3)-5 . From here solve for x

OpenStudy (anonymous):

he wants the inverse.. so @Princer_Jones i dont think your idea works..

OpenStudy (anonymous):

@magepker728 it works, read completely

OpenStudy (anonymous):

y=4(x+3)-5 implies y=4x+12-5 this implies y=4x+7 hence x=(y-7)/4 Now replacing y by x we have The inverse function is (x-7)/4 putting x=3 , the answer is -16

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

@bblock5 read the last post again

OpenStudy (calculusfunctions):

To determine algebraically the inverse of a function f(x) i). rewrite y = f(x) ii). interchange x and y iii). solve for y iv). write\[y =f ^{-1}(x)\]

OpenStudy (anonymous):

Ok thank you. If you guys dont mind, i have more questions

OpenStudy (calculusfunctions):

@Princer_Jones I've been observing how you simply give out the answers. Why don't you try explaining instead. How is anyone supposed to learn?

OpenStudy (anonymous):

Thank you @calculusfunctions

OpenStudy (anonymous):

@calculusfunctions I solve the problem too dont just post the final answer

OpenStudy (anonymous):

Which function below is the inverse of f(x)

OpenStudy (anonymous):